# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v2

@@ 2.2.2 The relationship between $L$, $C$ and $n$
The primary objective of parallel imaging (PI) techniques is to reduce imaging time, which is achieved by acquiring fewer data points per coil and taking multiple coil measurements at the same time. From compressive sensing theory [candes_romberg_tao_2006, recht_fazel_parrilo_2010], it is well known that the incoherence in the sensing matrix is crucial for the successful solution of the problems in (13) and (14). And it is well accepted in the literature on pMRI that increasing the number of channels $C$ improves the incoherence of the sensing matrix E.
@@
In this context, the works by Otazo et al. [otazo2009distributed, Otazo_Kim_Axel_Sodickson_2010] should be mentioned,where they numerically assess the sufficient conditions on the number of measurements per coil $L$ and the number of coils $C$ for solving the $\ell_{0}$ norm minimization problem in (12) or the $\ell_{1}$ norm minimization problem in (13) to recover an $n$-sparse signal. The authors conclude that in the known calibration setup, when the number of coils is large enough, the required number of measurements per coil is very close to the signal sparsity for both the noiseless and noisy cases. To illustrate the numerical findings, we reproduced the reconstruction results from [otazo2009distributed] by solving the $\ell_{0}$ minimization problem (12), using the orthogonal matching pursuit (OMP) algorithm. The results can be found in Figure 1.
@@
[…]
Note that this condition does not apply to multi-coil acquisition in pMRI, since all coils share the same undersampling pattern, albeit with distinct coil sensitivity profiles. Hence, one might expect a reduction in incoherence and a deviation from the theoretical recovery bounds discussed in [Duarte_Sarvotham_Baron_Wakin_Baraniuk]. Additionally, the measurement ensemble in the pMRI setting must adhere to additional physical constraints, which makes the analysis within the compressive sensing framework more complex. By combining the numerical experiments with the aforementioned theory for the known calibration case, one could reasonably expect that a sufficient condition for robust reconstruction would adopt a similar form to that of the best-case scenario for the more complex auto-calibration problem in (1).
@@ 2.2.3 Further discussions
Essentially, auto-calibration methods depend on data redundancy. In the literature on pMRI, other sources of redundancy have been utilized for auto-calibration, each under different assumptions. Data driven auto-calibration methods, such as GRAPPA [griswold_jakob_heidemann_nittka_jellus_wang_kiefer_haase_2002] and SPIRiT [Lustig_Pauly_2010] can be viewed as interpolation methods by estimating linear relationships within the fully sampled k-space (e.g., kernel calibration) and enforcing that relationship to synthesize data values in place of unacquired data (e.g., data reconstruction). More recently, a data-driven method SAKE [shin_larson_ohliger_elad_pauly_vigneron_lustig_2014] structures the multi-coil dataset into a new data matrix that is designed to have low-rankness property. Then it solves a matrix completion problem to fill in the missing data. Joint estimation techniques attempt to iteratively estimate both the coil sensitivities and image contents while imposing some smoothness constraints on the sensitivity profiles, some examples include [Ying_Sheng_2007, Uecker_Hohage_Block_Frahm_2008]. It is worth noting that the effectiveness of these methods depends on whether the coil and signal adhere to the assumptions made. As a result, they may produce varying quality of reconstructions for different datasets.
@@ 3.1 Lifting
Using the lifting approach [Candès_Eldar_Strohmer_Voroninski_2013, Ahmed_Recht_Romberg_2014, ling_strohmer_2015], we can transform the biconvex constraint (22) into the problem of recovering a block-sparse matrix from linear measurements.
@@
$\boldsymbol{\Phi}\coloneq\begin{bmatrix}(\mathbf{\bar{q}}_{1}\bigotimes\mathbf{b}_{1})^{*}\\ \vdots\\ \mathbf{\bar{q}}_{N}\bigotimes\mathbf{b}_{N})^{*}\end{bmatrix}\in\mathbb{C}^{N\times kN}$ (24)
@@
$\textbf{X}=\begin{bmatrix}\textbf{X}_{1}|\textbf{X}_{2}|\cdots|\textbf{X}_{C}\end{bmatrix}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}^{T})\end{bmatrix}=\textbf{z}\otimes\textbf{H}\in\mathbf{C}^{kN\times C},$
@@
where $\textbf{Y}\in\mathbb{C}^{L\times C}$, $\mathbf{F}_{\Omega}\in\mathbb{C}^{L\times N}$, $\mathbf{\Phi}\in\mathbb{C}^{N\times kN}$, $\textbf{X}\in\mathbb{C}^{kN\times 1}$. We refer to $\mathbf{A}=\mathbf{F}_{\Omega}\mathbf{\Phi}$ as the forward matrix or measurement ensemble. Now the problem is to find the matrix $\mathbf{X}$ satisfying the linear constraint described by (25). To that end we will exploit the block sparsity of X.
@@ 3.2 Block sparsity of $\mathbf{X}$
Moreover, X is at most rank-k if $C$ is large. Methods that exploit the low-rankness include [Akçakaya_Basha_Goddu_Goepfert_Kissinger_Tarokh_Manning_Nezafat_2011, trzasko_manduca_2011, Holme_Rosenzweig_Ong_Wilke_Lustig_Uecker_2019] and also the the ENLIVE formulation (17) introduced in Section 2.2. Naturally, one could minimize over a linear combination $\|\mathbf{\tilde{X}}\|_{1,2}+\lambda\|\mathbf{\tilde{X}}\|_{*},\lambda>0$ in an attempt to simultaneously capture the low-rankness and sparse characteristics of $\mathbf{\tilde{X}}$. However, as demonstrated in [Oymak_Jalali_Fazel_Eldar_Hassibi_2015], this approach is no more effective than using a single norm in terms of the number of measurements required. Theoretically, only recovering for a single coil case by minimizing $\|\cdot\|_{*}$ requires $L=O(N+k)$. Also $l_{1,2}$ minimization is more computationally efficient than the nuclear norm minimization problem. Hence, we choose to only minimize with respect to $\|\cdot\|_{1,2}$, which (as we will prove) may already yield a solution with sparse structure and small nuclear norm.
@@ 3.3 Optimization problem
Since it is well known that the $\ell_{1,2}$-norm promotes the block sparse structure in $\mathbf{X}$ [eldar_kuppinger_bolcskei_2010, eldar_mishali_2009, flinth_2017], we solve the following convex program
@@
Once we get a solution $\hat{\textbf{X}}$ to the $\ell_{1,2}$ minimization problem (27)(28), we calculate the average of the best rank-one estimations of its columns to retrieve the common underlying signal. Recall the columns of $\hat{\textbf{X}}=\begin{bmatrix}\hat{\textbf{x}}_{1}&\cdots\hat{\textbf{x}}_{C}\end{bmatrix}$ and the truth $\textbf{X}^{0}=\begin{bmatrix}\textbf{X}^{0}_{1}&\cdots\textbf{X}^{0}_{C}\end{bmatrix}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}^{T})\end{bmatrix}$. We use the following Lemma from [ling_strohmer_2015] to compute an average of the estimates from $\hat{\mathbf{X}}$ and to ensure that it is close to z.
@@
The Lemma is a direct consequence of combining Corollary 3.3 from [ling_strohmer_2015] and using norm equivalency. Also notice if for any $\alpha_{0}\neq 0$, $\alpha_{0}\textbf{u}_{0}$ and $1/\alpha_{0}\textbf{v}_{0}$ is also a pair of solutions. However, the scalar ambiguity is not important for many applications, including pMRI. The reconstructed signal is retrieved as $\hat{\textbf{x}}=\boldsymbol{\Psi}^{*}\hat{\textbf{v}}$
@@
A comprehensive review of the algorithms to solve problems of the form (29) is not a focus of this paper. We refer to a few common approaches such as the forward-backward splitting (FBS) [Combettes_Patrick_Wajs_Valerie_2005], ADMM [wang_yang_yin_zhang_2008] and FISTA [Beck_Teboulle_2009] which have found numerous applications in the context of pMRI reconstruction. In the simulations part of this paper, we use a FISTA-type iterative method to solve the unconstrained problem. Detailed iterative updates are provided in the Appendix 7.2.
@@ 4 Main result
Given any integer $k$ and $N$, we define $T_{j}\subset[kN]$ by $T_{j}=\{(j-1)k+1,\cdots,jk\},j\in[N]$. Given an index set $S=\{s_{1},s_{2},\cdots\}\subset[N]$, define $S_{j}=T_{s_{j}}$ for $j\in|S|$ and $T:=\cup_{j\in|S|}S_{j}$. With slight abuse of notation, we will not distinguish $\mathcal{P}_{S}$, $\mathcal{P}_{T}$ and $\mathcal{P}_{T\times[C]}$ when it is clear from the context. For example, for $\textbf{x}\in C^{N}$, $\textbf{x}\in C^{kN}$ and $\textbf{X}\in C^{kN\times C}$, $\mathcal{P}_{S}$ means the projection onto $S$, $T$ and $T\times[C]$ respectively. We will also sometimes write the projections $\mathcal{P}_{S}$ as subscripts, for example, $\textbf{x}_{S}$ or $\textbf{X}_{S}$, when it is convenient.
@@ 4.1 Assumptions and main theorem
    • Sampled from some basis of $\mathbb{C}^{k}$: Each column $\textbf{h}_{l},l\in[C]$ are chosen independently and uniformly at random with replacement from the columns of an orthonormal basis $\textbf{W}\in\mathbb{C}^{k\times k}$.
    • Complex spherical: Each spanning coefficients $\textbf{h}_{l}\in\mathbb{C}^{k},l\in[C]$ are chosen independently and uniformly at random from the complex sphere $S_{\mathbb{C}}^{k-1}$
We briefly introduced some common choices of B and $\boldsymbol{\Psi}$ in Section 2, including the smoothness promoting polynomial or sinosudal basis for B and DCT or wavelet transformation for $\boldsymbol{\Psi}$. Here, we further comment on our choice of assumptions (A1)-(A4). (A1) and (A3) are regarding the sensing matrix. (A3) is a standard assumption in compressive sensing theory concerning the random sampling scheme for k-space/Fourier space measurements. For (A1), B is chosen uniformly at random from all possible row permutations of a fixed matrix $\textbf{B}_{0}$. The orthogonality of the columns of B is not affected by row permutations since $\textbf{B}_{i}^{*}\textbf{B}_{j}=\sigma(\textbf{B}_{i})^{*}\sigma(\textbf{B}_{j})$. With B being a matrix with randomly permuted rows, for any fixed h, the coil sensitivity $\mathbf{s}=\textbf{B}\textbf{h}$ will be randomly permuted copies of some underlying $\textbf{s}_{0}=\textbf{B}_{0}\textbf{h}$. On the other hand, if B were deterministic, any zero entries in $\textbf{s}=\textbf{B}\textbf{h}$ will make it impossible to recover uniquely the values of the image z at those locations. The randomness in B helps in providing an average case analysis for each coil. Assumption (A4) on H is critical for developing an average case analysis across coils to explain the improved results when more channels are included.
@@
Before moving on to stating the theorem, let us discuss the reason for imposing such a probability model on H. Consider the worst case scenario where all the $\textbf{h}_{i}$ are identical, then the measurements $\textbf{Y}=\textbf{A}\textbf{X}$ consist of identical columns. By including multiple channels, we are not gaining more information on X regardless of the number of coil used. The following proposition gives a formal statement of the worst case analysis. The proof follows exactly from Proposition 4.1 in [Eldar_Rauhut_2010].
@@
(Worst case analysis) Suppose there exist a $\textbf{X}_{1}\in\mathbb{C}^{kN\times 1}$ that the $\ell_{1,2}$ minimization fails to receiver from $\textbf{Y}_{1}=\textbf{A}\textbf{X}_{1}$. Then the $\ell_{1,2}$ minimization fails to recover X = $\begin{bmatrix}\textbf{X}_{1}&\textbf{X}_{1}&\cdots\textbf{X}_{1}\end{bmatrix}\in\mathbb{C}^{kN\times C}$ from $[\textbf{Y}_{1},\cdots,\textbf{Y}_{1}]=\textbf{A}\textbf{X}$.
@@
with probability $1-4N^{-\alpha}$ if
@@
Here, $\mu^{B}_{max}:=\max\limits_{i,j}\sqrt{L}|\mathbf{B}_{i,j}|$, $\mu^{\Psi}_{max}:=\max\limits_{i,j}\sqrt{N}|{\boldsymbol{\Psi}}_{i,j}|$, which measure the incoherence of the rows of B and of $\boldsymbol{\Psi}$, respectively. The constant $C_{\alpha}$ grows linearly with respect to $\alpha$.
@@
A detailed bound for the constants in the theorem can be found in the proofs in Section 6. Next, we use Lemma 1 to get rid of the $\sqrt{C}$ scaling in the error bound in (32).
@@
Remark: The result provides a sufficient condition for stable and robust recovery in the presence of noise and sparse deficiency in the signal. It is important to emphasize that if the signal is truly sparse and there is no noise present, perfect reconstruction is achieved (i.e., the error is zero). Hence, increasing the number of the coils does not provide any additional benefit when the number of measurements is sufficiently large. Furthermore, to understand the strength of the statement, note that if $S$ is the true support of an n-sparse $\textbf{z}_{0}$, $\|\mathcal{P}_{S^{c}}\textbf{X}_{0}\|_{1,2}=0$, then the number of required measurements $L$ is of the order of the sparsity level of $\textbf{X}_{0}$, up to log factors (i.e., $kn={{\|P_{S}(\textbf{X}_{0})\|}_{0}}$). For the constants in the bound of (C1), it will be meaningful when the magnitudes of the entries in $\mathbf{B}$ and $\boldsymbol{\Psi}$ do not vary too much. Optimally, when the columns of $\mathbf{B}$ and $\boldsymbol{\Psi}$ are chosen from the DFT matrices, we optimally have $\mu^{B}_{max}=1$, $\mu^{\Psi}_{max}=1$.
@@
Ideally, we would like to derive a proof for the setup when both B and $\mathbf{\Psi}$ are fixed, and the randomness comes from the support set $S$ of z alone. The result would be analogous to the conditioning of the random sub-matrices of a fixed sampling matrix, e.g. see Chapter 14 of [foucart_rauhut_2015]. Yet, we encountered some (at least to us) unsolvable difficulties in finding a realistic bound for the incoherence measure between the sub-blocks of the sampling matrix A. Nevertheless, we still provide a formal statement of the bound and some discussions in the Appendix 7.1.
@@
The proof of the main theorem follows a well-established proof outline of compressive sensing [foucart_rauhut_2015]. First, we refer to a sufficient condition from [flinth_2017] to ensure a robust and stable recovery. Subsequently, we analyze the circumstances under which these conditions can be met. We show in Lemma 8 that under assumptions (A1)-(A4), the sufficient conditions for recovery are satisfied with a higher probability as the number of coils increases. This means that, in practice, parallel imaging is likely to perform better than single imaging recovery.

### v3

@@ 2.2.2 The relationship between $L$, $C$ and $n$
The primary objective of parallel imaging (PI) techniques is to reduce imaging time, which is achieved by acquiring fewer data points per coil and taking multiple coil measurements at the same time. From compressive sensing theory [48, 49], it is well known that the incoherence in the sensing matrix is crucial for the successful solution of the problems in (13) and (14). And it is well accepted in the literature on pMRI that increasing the number of channels $C$ improves the incoherence of the sensing matrix E.
@@
In this context, the works by Otazo et al. [50, 51] should be mentioned,where they numerically assess the sufficient conditions on the number of measurements per coil $L$ and the number of coils $C$ for solving the $\ell_{0}$ norm minimization problem in (12) or the $\ell_{1}$ norm minimization problem in (13) to recover an $n$-sparse signal. The authors conclude that in the known calibration setup, when the number of coils is large enough, the required number of measurements per coil is very close to the signal sparsity for both the noiseless and noisy cases. To illustrate the numerical findings, we reproduced the reconstruction results from [50] by solving the $\ell_{0}$ minimization problem (12), using the orthogonal matching pursuit (OMP) algorithm. The results can be found in Figure 1.
@@
Note that a key difference between the above problem and the pMRI forward model is that the sensors use different random undersampling patterns. The authors in [52] have theoretically shown that, in such a distributed compressed sensing model with a large number of sensors, the required number of samples per sensor approaches the sparsity level, as solved by a greedy OMP-type algorithm. Recall that compressed sensing with $\ell_{1}$-norm minimization (i.e. solving equation (13) with a single coil) typically requires $L$ to be three to five times the number of sparse coefficients $n$ (i.e., 335n5n where x is n-sparse) to achieve accurate reconstructions [30, 53]. Hence, a large number of sensors provides a constant improvement in the number of measurements per sensor, up to the sparsity level of the signal x.
Note that this condition does not apply to multi-coil acquisition in pMRI, since all coils share the same undersampling pattern, albeit with distinct coil sensitivity profiles. Hence, one might expect a reduction in incoherence and a deviation from the theoretical recovery bounds discussed in [52]. Additionally, the measurement ensemble in the pMRI setting must adhere to additional physical constraints, which makes the analysis within the compressive sensing framework more complex. By combining the numerical experiments with the aforementioned theory for the known calibration case, one could reasonably expect that a sufficient condition for robust reconstruction would adopt a similar form to that of the best-case scenario for the more complex auto-calibration problem in (1).
@@ 2.2.3 Further discussions
Essentially, auto-calibration methods depend on data redundancy. In the literature on pMRI, other sources of redundancy have been utilized for auto-calibration, each under different assumptions. Data driven auto-calibration methods, such as GRAPPA [44] and SPIRiT [54] can be viewed as interpolation methods by estimating linear relationships within the fully sampled k-space (e.g., kernel calibration) and enforcing that relationship to synthesize data values in place of unacquired data (e.g., data reconstruction). More recently, a data-driven method SAKE [31] structures the multi-coil dataset into a new data matrix that is designed to have low-rankness property. Then it solves a matrix completion problem to fill in the missing data. Joint estimation techniques attempt to iteratively estimate both the coil sensitivities and image contents while imposing some smoothness constraints on the sensitivity profiles, some examples include [6, 8]. It is worth noting that the effectiveness of these methods depends on whether the coil and signal adhere to the assumptions made. As a result, they may produce varying quality of reconstructions for different datasets.
@@ 3.1 Lifting
Using the lifting approach [2, 3, 1], we can transform the biconvex constraint (22) into the problem of recovering a block-sparse matrix from linear measurements.
@@
$\boldsymbol{\Phi}\coloneq\begin{bmatrix}(\mathbf{\bar{q}}_{1}\bigotimes\mathbf{b}_{1})^{*}\\ \vdots\\ (\mathbf{\bar{q}}_{N}\bigotimes\mathbf{b}_{N})^{*}\end{bmatrix}\in\mathbb{C}^{N\times kN}.$ (24)
@@
$\textbf{X}=\begin{bmatrix}\textbf{X}_{1}|\textbf{X}_{2}|\cdots|\textbf{X}_{C}\end{bmatrix}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}^{T})\end{bmatrix}=\textbf{z}\otimes\textbf{H}\in\mathbb{C}^{kN\times C},$
@@
where $\textbf{Y}\in\mathbb{C}^{L\times C}$, $\mathbf{F}_{\Omega}\in\mathbb{C}^{L\times N}$, $\mathbf{\Phi}\in\mathbb{C}^{N\times kN}$, $\textbf{X}\in\mathbb{C}^{kN\times C}$. We refer to $\mathbf{A}=\mathbf{F}_{\Omega}\mathbf{\Phi}$ as the forward matrix or measurement ensemble. Now the problem is to find the matrix $\mathbf{X}$ satisfying the linear constraint described by (25). To that end we will exploit the block sparsity of X.
@@ 3.2 Block sparsity of $\mathbf{X}$
Moreover, X is at most rank-k if $C$ is large. Methods that exploit the low-rankness include [55, 56, 47] and also the the ENLIVE formulation (17) introduced in Section 2.2. Naturally, one could minimize over a linear combination $\|\mathbf{\tilde{X}}\|_{1,2}+\lambda\|\mathbf{\tilde{X}}\|_{*},\lambda>0$ in an attempt to simultaneously capture the low-rankness and sparse characteristics of $\mathbf{\tilde{X}}$. However, as demonstrated in [27], this approach is no more effective than using a single norm in terms of the number of measurements required. Theoretically, only recovering for a single coil case by minimizing $\|\cdot\|_{*}$ requires $L=O(N+k)$. Also $l_{1,2}$ minimization is more computationally efficient than the nuclear norm minimization problem. Hence, we choose to only minimize with respect to $\|\cdot\|_{1,2}$, which (as we will prove) may already yield a solution with sparse structure and small nuclear norm.
@@ 3.3 Optimization problem
Since it is well known that the $\ell_{1,2}$-norm promotes the block sparse structure in $\mathbf{X}$ [57, 58, 4], we solve the following convex program
@@
Once we get a solution $\hat{\textbf{X}}$ to the $\ell_{1,2}$ minimization problem (27)(28), we calculate the average of the best rank-one estimations of its columns to retrieve the common underlying signal. Recall the columns of $\hat{\textbf{X}}=\begin{bmatrix}\hat{\textbf{x}}_{1}&\cdots\hat{\textbf{x}}_{C}\end{bmatrix}$ and the truth $\textbf{X}^{0}=\begin{bmatrix}\textbf{X}^{0}_{1}&\cdots\textbf{X}^{0}_{C}\end{bmatrix}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}^{T})\end{bmatrix}$. We use the following Lemma from [1] to compute an average of the estimates from $\hat{\mathbf{X}}$ and to ensure that it is close to z.
@@
The Lemma is a direct consequence of combining Corollary 3.3 from [1] and using norm equivalency. Also notice if for any $\alpha_{0}\neq 0$, $\alpha_{0}\textbf{u}_{0}$ and $1/\alpha_{0}\textbf{v}_{0}$ is also a pair of solutions. However, the scalar ambiguity is not important for many applications, including pMRI. The reconstructed signal is retrieved as $\hat{\textbf{x}}=\boldsymbol{\Psi}^{*}\hat{\textbf{v}}$ .
@@
A comprehensive review of the algorithms to solve problems of the form (29) is not a focus of this paper. We refer to a few common approaches such as the forward-backward splitting (FBS) [59], ADMM [60] and FISTA [61] which have found numerous applications in the context of pMRI reconstruction. In the simulations part of this paper, we use a FISTA-type iterative method to solve the unconstrained problem. Detailed iterative updates are provided in the Appendix 7.2.
@@ 4 Main result
Given any integer $k$ and $N$, we define $T_{j}\subset[kN]$ by $T_{j}=\{(j-1)k+1,\cdots,jk\},j\in[N]$. Given an index set $S=\{s_{1},s_{2},\cdots\}\subset[N]$, define $S_{j}=T_{s_{j}}$ for $j\in|S|$ and $T:=\cup_{j\in|S|}S_{j}$. With slight abuse of notation, we will not distinguish $\mathcal{P}_{S}$, $\mathcal{P}_{T}$ and $\mathcal{P}_{T\times[C]}$ when it is clear from the context. For example, for $\textbf{x}\in\mathbb{C}^{N}$, $\textbf{x}\in\mathbb{C}^{kN}$ and $\textbf{X}\in\mathbb{C}^{kN\times C}$, $\mathcal{P}_{S}$ means the projection onto $S$, $T$ and $T\times[C]$ respectively. We will also sometimes write the projections $\mathcal{P}_{S}$ as subscripts, for example, $\textbf{x}_{S}$ or $\textbf{X}_{S}$, when it is convenient.
@@ 4.1 Assumptions and main theorem
    • Sampled from some basis of $\mathbb{C}^{k}$: Each column $\textbf{h}_{l},l\in[C]$ is chosen independently and uniformly at random with replacement from the columns of an orthonormal basis $\textbf{W}\in\mathbb{C}^{k\times k}$.
    • Complex spherical: Each spanning coefficients $\textbf{h}_{l}\in\mathbb{C}^{k},l\in[C]$ is chosen independently and uniformly at random from the complex sphere $S_{\mathbb{C}}^{k-1}$
We briefly introduced some common choices of B and $\boldsymbol{\Psi}$ in Section 2, including the smoothness promoting polynomial or sinusoidal basis for B and DCT or wavelet transformation for $\boldsymbol{\Psi}$. Here, we further comment on our choice of assumptions (A1)-(A4). (A1) and (A3) are regarding the sensing matrix. (A3) is a standard assumption in compressive sensing theory concerning the random sampling scheme for k-space/Fourier space measurements. For (A1), B is chosen uniformly at random from all possible row permutations of a fixed matrix $\textbf{B}_{0}$. The orthogonality of the columns of B is not affected by row permutations since $\textbf{B}_{i}^{*}\textbf{B}_{j}=\sigma(\textbf{B}_{i})^{*}\sigma(\textbf{B}_{j})$. With B being a matrix with randomly permuted rows, for any fixed h, the coil sensitivity $\mathbf{s}=\textbf{B}\textbf{h}$ will be randomly permuted copies of some underlying $\textbf{s}_{0}=\textbf{B}_{0}\textbf{h}$. The randomness in B helps in providing an average case analysis for each coil. Practically, if a coil $\textbf{s}=\textbf{B}\textbf{h}$ were deterministic, any zero entries in s (e.g. dead pixels or inactive responses) would result in a loss of information about the object of interest z at those corresponding locations. Assumption (A4) on H is critical for developing an average case analysis across coils to explain the improved results when more channels are included.
@@
Before moving on to stating the theorem, let us discuss the reason for imposing such a probability model on H. Consider the worst case scenario where all the $\textbf{h}_{i}$ are identical, then the measurements $\textbf{Y}=\textbf{A}\textbf{X}$ consist of identical columns. By including multiple channels, we are not gaining more information on X regardless of the number of coil used. The following proposition gives a formal statement of the worst case analysis. The proof follows exactly from Proposition 4.1 in [62].
@@
(Worst case analysis) Suppose there exist a $\textbf{X}_{1}\in\mathbb{C}^{kN\times 1}$ that the $\ell_{1,2}$ minimization fails to receiver from $\textbf{Y}_{1}=\textbf{A}\textbf{X}_{1}$. Then the $\ell_{1,2}$ minimization fails to recover X = $\begin{bmatrix}\textbf{X}_{1}&\textbf{X}_{1}&\cdots\textbf{X}_{1}\end{bmatrix}\in\mathbb{C}^{kN\times C}$ from $\textbf{Y}_{1},\cdots,\textbf{Y}_{1}]=\textbf{A}\textbf{X}$.
@@
with probability $1-4N^{-\alpha}$ if the following conditions are met,
@@
Here, $\mu^{B}_{max}:=\max\limits_{i,j}\sqrt{L}|\mathbf{B}_{i,j}|$, $\mu^{\Psi}_{max}:=\max\limits_{i,j}\sqrt{N}|{\boldsymbol{\Psi}}_{i,j}|$, which measure the incoherence of the rows of B and of $\boldsymbol{\Psi}$, respectively. Note that this condition includes many useful examples–for instance, when $\boldsymbol{\Psi}$ is the unitary DFT matrix, the constant $\mu^{\Psi}_{max}$ achieves the optimal value of one  [^2] A relaxed version allowing the constants to grow with the logarithmic scale of $N$ is also possible. The constant $C_{\alpha}$ grows linearly with respect to $\alpha$.
@@
A detailed bound for the constants in the theorem can be found in the proofs in Section 6. Next, we use Lemma 1 to get back the signal of interest from the overall lifted matrix $\mathbf{X}$, which also eliminates the $\sqrt{C}$ scaling in the error bound in (32).
@@
Remark: The result provides a sufficient condition for stable and robust recovery in the presence of noise and sparse deficiency in the signal, requiring both the number of the measurements per coil and the total number of coils to be sufficiently large for the boundary case. It is important to emphasize that if the signal is truly sparse and there is no noise present, perfect reconstruction is achieved (i.e., the error is zero). Our theoretical bound exhibits the same form as the empirical findings in [51] (even in their non-convex setup) and is validated by our numerical tests. Furthermore, it also implies that once $L$ and $C$ are both sufficiently large, increasing the number of coils does not yield any additional benefit in reducing the number of measurements per coil or the reconstruction quality. In contrast, as we will show in Lemma 8, for small $C$, increasing the number of coils improves the probability of recovery. Furthermore, to understand the strength of the statement, note that if $S$ is the true support of an $n$-sparse $\textbf{z}_{0}$, $\|\mathcal{P}_{S^{c}}\textbf{X}_{0}\|_{1,2}=0$, then the number of required measurements $L$ is of the order of the sparsity level of $\textbf{X}_{0}$, up to log factors (i.e., $kn={{\|P_{S}(\textbf{X}_{0})\|}_{0}}$). For the constants in the bound of (C1), it will be meaningful when the magnitudes of the entries in $\mathbf{B}$ and $\boldsymbol{\Psi}$ do not vary too much. Optimally, when the columns of $\mathbf{B}$ and $\boldsymbol{\Psi}$ are chosen from the DFT matrices, we optimally have $\mu^{B}_{max}=1$, $\mu^{\Psi}_{max}=1$.
@@
Ideally, we would like to derive a proof for the setup when both B and $\mathbf{\Psi}$ are fixed, and the randomness comes from the support set $S$ of z alone. The result would be analogous to the conditioning of the random sub-matrices of a fixed sampling matrix, e.g. see Chapter 14 of [21]. Yet, we encountered some (at least to us) unsolvable difficulties in finding a realistic bound for the incoherence measure between the sub-blocks of the sampling matrix A. Nevertheless, we still provide a formal statement of the bound and some discussions in the Appendix 7.1.
@@
The proof of the main theorem follows a well-established proof outline of compressive sensing [21]. First, we refer to a sufficient condition from [4] to ensure a robust and stable recovery. Subsequently, we analyze the circumstances under which these conditions can be met. We show in Lemma 8 that under assumptions (A1)-(A4), the sufficient conditions for recovery are satisfied with a higher probability as the number of coils increases. This means that, in practice, parallel imaging is likely to perform better than single imaging recovery.